Problem 9. Variance Analysis Bronfenbrenner Co. uses a standard cost system for its single product in which variable overhead is applied on the basis of direct labor hours. The following information is given: Standard costs per unit: Raw materials (1.5 grams x S16 per gram) Direct labor (0.75 hours x S8 per hour) S6.00 Variable overhead (0.75 hours × $3 per $24.00 S2.25 hour) Actual experience for current year Units produced Purchases of raw materials (21,000 grams x S17 per gram) 22 ,400 units $357,000 33,400 grams Raw materials used Direct labor (16,750 hours x $8 per hour) Varia Required Compute the following variances for raw materials, direct labor, and variable overhead, assuming that the price variance for materials is recognized at point of purchase: a. Direct materials price variance. $134,000 ble overhead cost incurred S48,575 b. Direct materials quantity variance. c. Labor rate variance d. Labor efficiency variance. e. Variable overhead rate variance. f. Variable overheadefficiency variance.
Answer #1
Req a:
Material Price variance:
Std price per unit: 16 per unit
Actual price per unit: 17 per unit
Actual Qty purchased: 21000 units
Material price variance: Actual Qty purchased (Std price-Actual price)
21000 (16-17) = 21000 Unfav
Req b:
Std price:16
Actual qty used: 33400
Std Qty allowed (22400*1.50) = 33600
Material Qty variance: Std price (Std qty-Actual qty used)
16 (33600-33400) = 3200 Fav
Req c:
Std rate per hour; 8
Actual rate per hour: 8
Actual hours worked: 16750
Labour rate variance: Actual hours (Std rate-Actual rate)
16750 (8-8) = Nil
Req d:
Std rate: 8
Actual Hours: 16750
Std hours (22400*0.75) = 16800
Labourn efficiency variance = Std rate (Std hours-Actual hours)
8 ( 16800 -16750 )= 400 fav
Req e:
Std variable OH rate peer hour: 3
Actual rate per hour (48575/16750) = 2.90 per hour
Actual hours: 16750
Vvariable OH rate variance: Actual hours (Std rate - Aactual rate)
16750 ( 3.00-2.90) = 1675 Fav
Req f:
Std rate: 3.00
Actual hours: 16750
Std hours allowed: 16800
Variable OH efficiency variance: Std rate (Std hours-Actual hours)
3.00 ( 16800-16750) = 150 fav